OpenStudy (diamond77108):

PLZZ HELP WILL MEDAL AND FAN!!!! Which statement about the ordered pairs (3, −8) and (4, 4) is true for the equation 3x−y/4=11? (3, −8) is a solution to the equation. (4,  4) is a solution to the equation. Both ordered pairs are solutions. Neither ordered pair is a solution.

OpenStudy (diamond77108):

\[3x - \frac{ y }{ 4 }= 11\]

OpenStudy (mecharv):

Clearly, both ordered pairs are a solution.

OpenStudy (diamond77108):

thx

OpenStudy (mecharv):

First one (3,-8) being 3(3)-(-8/4)=9+2=11 Second one being (4,4), 3(4)-(4/4)=12-1=11

OpenStudy (diamond77108):

can you help with one more

OpenStudy (mecharv):

I am not going to do your entire assignment. Please understand and proceed on your own.

OpenStudy (diamond77108):

alr thx anyway